OpenStudy (anonymous):

3x - 5y = 25 x = y + 7

OpenStudy (anonymous):

ummm so do u need to use subsitution or elimination?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

ok that is easy

OpenStudy (anonymous):

3(y+7)-5y=25

OpenStudy (anonymous):

Can u solve from there?

OpenStudy (anonymous):

can you show me the next step? then I think I can from there... I understand a little better

OpenStudy (anonymous):

so we open up the brackets and gather like terms 3y+21-5y=25 21-2y=25 -2y=3 y=(-3/2) Now plug this into the second equation to find out what x equals

OpenStudy (anonymous):

Then to check your work plug in the x and y into both equations and if the answer is correct then you know u got the correct answer

OpenStudy (anonymous):

okay! Thank you so much. I think I got it now.

OpenStudy (anonymous):

alrighty :D